(46) Science and medicine. Celsius temperature readings can be converted to Fahrenheit readings using the formula F=9/5C + 32. What is the Fahrenheit temperature that corresponds to eachof the following Celsius temperatures: -10, 0,15,100?
(48)Business and finance. When x number of units are sold, the price of each unit(in dollars) is given by p= -x/2 +75. Find the unit price when the following quantities are sold:2,7,9,11.
(64) Business and finance. The revenue for a sandwich shop is directly proportional to its advertising budget.When the owner spent $2000 a month on advertising. The revenue was $120,000. If the revenue is now $180,000. how much is the owner spending on advertising?

Answer by checkley71(8403)

F=9/5(-10)+32
F=-90/5+32
F=-18+32
F=14
----------------------------------
F=9/5(0)+32
F=32
-----------------------------------
F=9/5(15)+32
F=135/5+32
F=27+32
F=59
---------------------------------
F=9/5(100)+32
F=900/5+32
F=180+32
F=212
#######################################################
MERRY CHRISTMAS